mathslover (mathslover):

Solve the following equation for the variable: \[x^2=4\]

mathslover (mathslover):

\[x=\pm 2\] ?

mathslover (mathslover):

@ash2326 @amistre64

OpenStudy (andriod09):

yes.

mathslover (mathslover):

what if : \(x=\sqrt{4}\)

OpenStudy (andriod09):

\[x ^{2}=+4\]

mathslover (mathslover):

I wonder that there is something I am forgetting.. @ash2326 can u remind me?

mathslover (mathslover):

x = sqrt4 and x^2 = 4

mathslover (mathslover):

isn't something like this? that has two diff. soln ?

OpenStudy (ash2326):

if we have \[x=\sqrt 4\] then x=2 if \[x^2=4\] \[x^2-4=0\] \[(x+4)(x-4)\] \[x=\pm 2\]

OpenStudy (ash2326):

oops (x+2)(x-2)

mathslover (mathslover):

Exactly but why is it so that x = \sqrt{4} then we have x = 2 ? and not x = -2..

OpenStudy (ash2326):

square root is defined like this. Its range is positive real no.s only

mathslover (mathslover):

any example?

ganeshie8 (ganeshie8):

\(x^2 = 4\) => \(x = \pm\sqrt{4}\)

OpenStudy (ash2326):

your question is an example of this \[y=\sqrt {x}\] domain = \(x\ge 0\) range=\(y \ge=0\)

mathslover (mathslover):

Oh gotcha... Thanks @ash2326 and @ganeshie8 (:))

OpenStudy (ash2326):

welcome @mathslover :)

OpenStudy (anonymous):

it's like this\[x^2=4\] then take mode of both sides \[|x^2|=4\]--->\[|x|^2=4\]---->\[|x|=2\]----->\[x=+2,-2\]